OpenStudy (anonymous):

A toy factory makes a red toy, a green toy, and a blue toy. In a lot of 2,500 toys, 500 were red, 1,500 were green, and 500 were blue. The toy factory has a defect rate of 4%, regardless of the color. If a toy is chosen at random, what is the probability it is blue or red, and defective?

OpenStudy (anonymous):

Number of defective blue toys = 0.04 * 500 = 20 Number of defective red toys = 0.04 * 500 = 20 40 defective blue and red toys Probability = 40 / 2500...which simplifies to 2 / 125 if you want your answer as a fraction or 1.6% if you want a percentage or 0.016 if you want a decimal

OpenStudy (anonymous):

Ok...think I got rid of the typos. Should be right.